The following transactions pertain to 2018, the first-year operations of Franklin Company. All inventory was started and completed during 2018. Assume that all transactions are cash transactions. Acquired $4,600 cash by issuing common stock. Paid $640 for materials used to produce inventory. Paid $1,910 to production workers. Paid $643 rental fee for production equipment. Paid $110 to administrative employees. Paid $105 rental fee for administrative office equipment. Produced 310 units of inventory of which 230 units were sold at a price of $12 each. Required Prepare a balance sheet in accordance with GAAP.
Answer #1
Balance sheet
Assets
Cash 3942
Inventory80*10.3 824
Total assets 4766
Equity
Retained earnings 166
Common stock 4600
Total equity 4766

Working notes:

Cost per unit = Total expense/Total production

Total expense = 640+1910+643 = 3193

Cost per unit = 3193/310 = 10.3

Income statement
Sales revenue(230*12) 2760
Less: Cost of goods sold(230*10.3) 2369
Gross margin 391
Less: Admin expense
         Admin employees 110
         Admin rental 115
Net income 166